OpenStudy (anonymous):

On a buying trip in Los Angeles, Rosaria Perez ordered 120 pieces of jewelry: a number of bracelets at $4 each and a number of necklaces at $13 each. She wrote a check for $1110 to pay for the order. How many bracelets and how many necklaces did Rosaria purchase? On a buying trip in Los Angeles, Rosaria Perez ordered 120 pieces of jewelry: a number of bracelets at $4 each and a number of necklaces at $13 each. She wrote a check for $1110 to pay for the order. How many bracelets and how many necklaces did Rosaria purchase? @Mathematics

OpenStudy (anonymous):

4x + 13 (120-x) = 1110 Braclets = 50 Necklaces = 70

OpenStudy (anonymous):

Let number of Bracelets = x then number of necklaces = (120 - x) cost of x bracelets = x*4 = 4x cost of (120-x) necklaces = (120-x)*13 = (1560 - 13x) ATQ, 4x + (1560 - 13x) = 1110 4x - 13x = 1110 - 1560 -9x = - 450 x = -450/-9 x = 50 so number of bracelets = x = 50 number of necklaces = 120 - 50 = 70